### v2

# TextCrafter: Accurately Rendering Multiple Texts in Complex Visual Scenes

Abstract
This paper explores the task of Complex Visual Text Generation (CVTG), which centers on generating intricate textual content distributed across diverse regions within visual images. In CVTG, image generation models often rendering distorted and blurred visual text or missing some visual text. To tackle these challenges, we propose TextCrafter, a novel multi-visual text rendering method. TextCrafter employs a progressive strategy to decompose complex visual text into distinct components while ensuring robust alignment between textual content and its visual carrier. Additionally, it incorporates a token focus enhancement mechanism to amplify the prominence of visual text during the generation process. TextCrafter effectively addresses key challenges in CVTG tasks, such as text confusion, omissions, and blurriness. Moreover, we present a new benchmark dataset, CVTG-2K, tailored to rigorously evaluate the performance of generative models on CVTG tasks. Extensive experiments demonstrate that our method surpasses state-of-the-art approaches.
Figure 1: TextCrafter enables precise multi-region visual text rendering, addressing the challenges of long, small-size,various numbers, symbols and styles in visual text generation. We illustrate the comparisons among TextCrafter with three state-of-the-art models, i.e., FLUX, TextDiffuser-2 and 3DIS.
1 Introduction
In recent years, diffusion models Ho et al. (2020); Podell et al. (2023); Rombach et al. (2022); Ramesh et al. (2021, 2022); Zhao et al. (2024); Dhariwal and Nichol (2021); Chen et al. (2023) have emerged as the forefront technology in visual image generation. Recent studies, represented by general-purpose text-to-image models such as Flux BlackForest (2024) and SD 3.5 Esser et al. (2024), have acquired the ability to render simple visual text through large-scale pretraining. However, when confronted with complex real-world visual text scenarios, these models often struggle with critical challenges such as text distortion, omission, and blurriness. These shortcomings significantly hinder their practical applicability in real-world use cases. Achieving precise, legible, and contextually accurate visual text rendering in complex scenes remains essential yet highly challenging.
Recent studies in multi-instance control Feng et al. (2022); Zhou et al. (2024); Chen et al. (2024b); Zhou et al. (2025) primarily focus on generating multiple instances within a single image simultaneously, striving to achieve precise regulation over factors such as quantity, position and attributes. However, a crucial yet often overlooked category in the visual domain is visual text. Text is an indispensable component of the real world, distinguished by its unparalleled complexity and delicacy compared to conventional objects, flora, or fauna. Unlike these entities, text possesses an exceptionally intricate internal structure, where even minor alterations can cause significant changes in visual appearance, potentially leading to misrecognition or complete illegibility. Existing multi-instance approaches, such as MIGC Zhou et al. (2024) and 3DISZhou et al. (2025), primarily focus on sentence-level information and interaction, offering a limited understanding of text at a finer granularity. Moreover, RPG Yang et al. (2024a) and RAG-Diffusion Chen et al. (2024b) adjust instance structures during attention fusion but fall short of the precision and stability needed for visual text rendering.
Recent studies on visual text rendering primarily focus on accurately generating a single text region. For instance, methods like AnyText Tuo et al. (2024), Glyph-byT5 Liu et al. (2025), SceneTextGen Zhangli et al. (2024), Diff-Text Zhang et al. (2024b), and TextDiffuser Chen et al. (2024a) often employ specialized modules to encode visual text representations and condition the diffusion model accordingly. However, these methods face two major limitations: 1) They rely on predefined rules to synthesize visual text training datasets. Due to the need for highly accurate text annotations, manual verification is often required, making the creation of complex visual text datasets both labor-intensive and costly. 2) Most approaches depend on fine-tuned text encoders (e.g., Glyph-by-T5) or trained conditional control encoders (e.g., AnyText, Diff-Text, TextDiffuser, SceneTextGen) to facilitate text generation. In complex multi-text generation, however, these approaches often cause interference among the control features of different targets, making it difficult to balance fine-grained local control with global consistency across multiple texts.
This paper explores the task of Complex Visual Text Generation (CVTG), which aims to generate intricate text content distributed across different regions. To tackle the challenges of CVTG, we propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes. TextCrafter enhances multi-text control, ensuring accurate text presentation and logical layout. Our framework comprises three key stages: 1) Instance Fusion: This stage strengthens the connection between visual text and its corresponding carrier, ensuring strong consistency between the text content and its surrounding environment. It effectively prevents text from appearing in incorrect positions. 2) Region Insulation: Leveraging the positional priors of the pre-trained DiT model Peebles and Xie (2023) (e.g., FLUX), this step initializes the layout information for each text instance while separating and denoising text prompts across different regions. This process prevents early interference between text areas, reducing confusion and the risk of content omission in multi-text scenarios. 3) Text Focus: An attention control mechanism is introduced to enhance the attention maps of visual text, refining the fidelity of text rendering. This stage ensures textual accuracy and addresses issues of blurriness, particularly in smaller text. We present some examples in Figure 1. Our contributions are summarized as follows:
  • We propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes, consisting of three key stages: instance fusion, region insulation, and token focus.
  • We construct a new CVTG-2K dataset, which encompasses a diverse range of visual text prompts varying in position, quantity, length, and attributes, serving as a robust benchmark for evaluating the performance of generation models in CVTG tasks.
  • We conduct extensive quantitative and qualitative experiments on the CVTG-2K benchmark. The results demonstrate that TextCrafter delivers exceptional performance, proving its superior effectiveness and robustness in tackling the CVTG task.
Figure 2: Overview of our TextCrafter. TextCrafter consists of three steps. (a) Instance Fusion: Strengthen the connection between visual text and its corresponding carrier. (b) Region Insulation: Leverage the positional priors of the pre-trained DiT model to initialize the layout information for each text instance while separating and denoising text prompts across different regions. (c) Text Focus: Enhance the attention maps of visual text, refing the fidelity of text rendering.
2 Related Works
Multi-instance Compositional Generation.
Attend-and-Excite Chefer et al. (2023) and BOX-Diffusion Xie et al. (2023) guide pre-trained diffusion models to align generated instances with prompts. GLIGEN Li et al. (2023), MIGC Zhou et al. (2024), and CreativeLayout Zhang et al. (2024a) enhance control by incorporating bounding box conditions into trainable layers. RPG Yang et al. (2024a) and RAG-Diffusion Chen et al. (2024b) break the generation process into regional tasks for compositional generation. However, these methods often overlook a critical element: visual text. As a key component of visual scenes, accurate text significantly enhances the realism of generated images.
Visual Text Generation.
DiffSTE Ji et al. (2023), Glyph-ByT5 Liu et al. (2025), and UDiffText Zhao and Lian (2025) use character-level encoders to incorporate word appearance into embeddings. TextDiffuser Chen et al. (2024a) and GlyphControl Yang et al. (2024b) generate text layout masks for injection into the latent space. Approaches like Diff-Text Zhang et al. (2024b) improve accuracy with attention map restrictions. GlyphDraw Ma et al. (2023) uses two models for text position prediction and rendering. TextDiffuser-2 Chen et al. (2025) leverages large language models for layout planning. However, these approaches often optimize text rendering for single region or on individual object, leading to poor generalization in complex visual text scenes.
Benchmark for Visual Text.
Earlier works like GlyphControl introduced SimpleBench and CreativeBench Yang et al. (2024b), but these fixed-template datasets offer limited diversity and contain only single-word scenes. MARIO-Eval Chen et al. (2024a) suffers from low-quality prompts and unclear semantics. DrawTextExt Creative Ma et al. (2023) collects prompts for generating natural scenes but still lack sufficient coverage. AnyText-benchmark Tuo et al. (2024) allows multi-word prompts but treats each word separately and appends them directly to the caption, limiting data distribution. Existing visual text benchmarks focus on text generation within a single object or location, falling short of capturing the complexity of real-world scenes. Thus, we propose a benchmark CVTG-2K to incorporate diverse positions, quantities, lengths, and attributes for comprehensively evaluate models in complex visual text scenarios.
Figure 3: Illustration of tokenizing the prompt “A sidewalk poster with ‘Register Now for IJCAI 2025’.” along with the attention map corresponding to each token. The use of preceding quotation marks can reinforce the relationship between text tokens and carrier tokens.
3 Methodology
3.1 Overview
Problem Definition.
In CVTG, the user provides a global prompt $P$ that contains multiple descriptions of visual texts $D=\{d_{1},d_{2},\dots,d_{n}\}$, where each description includes the main body of the visual text and other descriptors such as the position and attributes of the visual text. The main body of the visual text is defined as $VT=\{vt_{1},vt_{2},\dots,vt_{n}\}$, where each $vt_{i}$ corresponds to all of its descriptors $d_{i}$. The goal of the model is to generate an image based on the provided prompt $P$, where each visual text $vt_{i}$ is presented with its corresponding description $d_{i}$.
Challenges in CVTG.
When handing complex prompts, diffusion models encounter interference among visual texts. This interference leads to several degradation issues. $1$) Text confusion: Visual texts (e.g., visual text $vt_{i}$ and visual text $vt_{j}$) intertwine, leading to the generation of erroneous character or word. $2$) Text omission: Only visual texts related to description $d_{i}$ are generated, neglecting the visual text $vt_{j}$ in description $d_{j}$. $3$) Text blurriness: Visual text in description $d_{i}$ is small in scale and fails to draw sufficient attention, causing $vt_{i}$ to appear blurred in generated image.
Motivation.
Current text-to-image models excel at handing individual visual text in visual generation. By breaking down CVTG into simpler sub-tasks, overall complexity is reduced. Solving these sub-tasks separately and combining their solutions effectively accomplishes the CVTG task. This approach directs the model to progressively shift focus from the entire task to individual sub-tasks and then to specific visual texts, enabling precise rendering. Based on this coarse-to-fine strategy, we present TextCrafter, as shown in Figure 2, with detailed explanations provided in the following sections.
3.2 Instance Fusion
It is crucial to prevent visual texts from floating in the image and ensure proper positioning. We propose an instance fusion strategy that integrates the visual text’s content with its spatial carrier. Rather than using the text’s own embedding, we leverage embedding of the preceding quotation mark to establish spatial relationships. As illustrated in Figure 3, the preceding quotation mark in the attention map corresponds to the visual text it governs, indicating that it encapsulates complete information about the text. According to token additivity Hu et al. (2024), we incorporate the quotation mark’s embedding into the carrier’s embedding via weighted fusion to ensure precise alignment, where $\lambda$ controls the proportion of the addition. . Given a prompt $P$, after being processed by the T5 Raffel et al. (2020) text encoder $\tau_{\xi}$, the condition is expressed as $C=\tau_{\xi}(P)$. Each row of $C$ represents the token embedding corresponding to each token processed by T5.
Figure 4: For a pre-trained DiT model, only a few denoising steps are required to approximate the layout of the image and the relative positions of the main subjects. After 8 denoising steps, the layout closely resembles that of a full 50-step process, with subsequent steps primarily refining image details.
Benchmark Number Word Char Attribute Region
CreativeBench 400 1.00 7.29 × ×
MARIOEval 5414 2.92 15.47 × ×
DrawTextExt 220 3.75 17.01 × ×
AnyText-benchmark 1000 4.18 21.84 × ×
CVTG-2K 2000 8.10 39.47
$\checkmark$
size/color/font
$\checkmark$
2/3/4/5
Table 1: Comparison of CVTG-2K with existing public visual text Benchmarks in terms of Number of Samples(Number), Average Word Count (Word), Average Character Count (Char), Attribute, and number of Regions (Region).
Algorithm 1 TextCrafter image generation
Input: The encoded embedding of $P$ and $D$ is denoted as $c$ and $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$
Parameter: the proportion of the addition $\lambda$, the $r$ steps of region insulation, the pre-generation steps $\tau$
Output: A generated latent $x_{0}$
1: {Step 1 Instance Fusion}
2:$c=\text{InstanceFusion}(c,\lambda)$
3: {Step 2 Region Insulation}
4:$z_{t}\sim\mathcal{N}(0,1)$
5:for $t=\tau,\tau-1,\dots,1$ do
6:    $z_{t-1},A_{t}\leftarrow\text{DiT}(z_{t},c,t)$
7:end for
8:for $i=1,2,\dots,n$ do
9:    $p_{i}\leftarrow(0,0)$
10:    $p_{\text{max},i}\leftarrow\arg\max_{p_{i}}A_{t}(p_{i},i)$
11:end for
12:$bbx_{1},bbx_{2},\dots,bbx_{n}\leftarrow\text{MILP}(p_{\text{max},1},p_{\text{max},2},\dots,p_{\text{max},n})$
13:$x_{T}\sim\mathcal{N}(0,1)$
14:for $i=1,2,\dots,n$ do
15:    $\hat{x}_{t}^{i}\leftarrow\text{Initialize}(\mathcal{N}(0,1),bbx_{i})$
16:end for
17:for $t=T,T-1,\dots,T-r+1$ do
18:    $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$
19:    for $i=1,2,\dots,n$ do
20:       $\hat{x}_{t-1}^{i}\leftarrow\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$
21:       $x_{t-1}\leftarrow\text{Reinsert}(x_{t-1},\hat{x}_{t-1}^{i},bbx_{i})$
22:    end for
23:end for
24: {Step 3 Text Focus}
25:for $t=T-r,T-r-1,\dots,1$ do
26:    for $AttentionLayer_{i}\in\epsilon_{\theta}$ do
27:       $M_{t}=Q_{t}^{i}\times K_{t}^{i}$
28:       $M_{t}\leftarrow\text{Focus}(M_{t})$
29:       $Attention_{t}^{i}\leftarrow\text{softmax}(M_{t}/\sqrt{d})\times V_{i}$
30:    end for
31:    $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$
32:end for
33:return $x_{0}$
Word Accuracy↑
Model 2 regions 3 regions 4 regions 5 regions average NED↑ CLIPScore↑
SD3.5 Large (ICML 2024) 0.7293 0.6825 0.6574 0.5940 0.6548 0.8470 0.7797
FLUX.1 dev (2024) 0.6089 0.5531 0.4661 0.4316 0.4965 0.6879 0.7401
AnyText (ICLR 2024) 0.0513 0.1739 0.1948 0.2249 0.1804 0.4675 0.7432
TextDiffuser-2 (ECCV 2024) 0.5322 0.3255 0.1787 0.0809 0.2326 0.4353 0.6765
RAG-Diffusion (Arxiv 2024) 0.4388 0.3316 0.2116 0.1910 0.2648 0.4498 0.7797
3DIS (ICLR 2025) 0.4495 0.3959 0.3880 0.3303 0.3813 0.6505 0.7767
TextCrafter (Ours) 0.7628 0.7628 0.7406 0.6977 0.7370 0.8679 0.7868
Table 2: Quantitative results on the proposed CVTG-2K. The results highlight TextCrafter’s superiority across three metrics.
3.3 Region Insulation
Pre-generation.
In CVTG, multiple texts often influence each other’s glyphs and layout Zheng et al. (2023). To simplify, we decouple them into independent instances. Specifically, we assign a rectangular bounding box to each visual text to represent its layout, initialized based on the positional preferences of the pre-trained DiT model. This allows us to effectively isolate the texts, preventing interference during generation. Instead of manually specifying layouts or relying on large language models, we leverage the model’s learned positional knowledge Tumanyan et al. (2023), as enforcing a fixed layout may compromise quality. As shown in Figure 4, during the early denoising steps, we capture attention maps and identify the points of maximum attention for each visual text. This pre-generation process guides layout initialization. Assume that the attention map obtained during this process is $A_{t}$, where $t$ represents the current step, and $p$ stands for pixel:
$p_{\text{max}}=\arg\max_{p}A_{t}(p).$ (1)
Layout Optimizer.
Next, the rectangular bounding boxes for each visual text are optimized using Mixed-Integer Linear Programming (MILP) to minimize the Manhattan distance between the bounding box center and the attention point. Let $p_{\text{max},i}$ be the maximum attention point of $vt_{i}$ and $c_{i}$ the rectangular bounding box center. The optimization objective is:
$\min\sum_{i}D_{\text{Manhattan}}(p_{\text{max},i},c_{i}).$ (2)
Simultaneously incorporate a set of constraints to ensure that the layout of each visual text is reasonable and non-overlapping. The solution to this optimization problem can be represented as assigning a bounding box $bbx_{i}$ $(m_{\text{offset}},n_{\text{offset}},m_{\text{scale}},n_{\text{scale}})$ to each description. Then, we encode the prompt $P$ and the corresponding descriptions of visual text $D=\{d_{1},d_{2},\dots,d_{n}\}$, obtaining values for $c$ and $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. $c$ needs to be processed through instance fusion introduced in Section 3.2. The latent $\hat{x}_{i}$ is conditioned on $\hat{c}_{i}$, and the initial latent $x$ is conditioned on the complete $c$. When $t>T-r$ (i.e., during the first $r$ steps of the denoising process), denoising follows the formula:
$x_{t-1}$ $=x_{t}-\epsilon_{\theta}(x_{t},c)$
$\hat{x}_{t-1}^{i}=\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i}),$ (3)
where $\epsilon_{\theta}$ represents the noise predictor. After each denoising step, $\hat{x}_{t}^{i}$ is reinserted into the corresponding region in $x_{t}$.
3.4 Text Focus
Region insulation creates distinct visual boundaries between visual texts. To enhance harmony, regions are merged into a single latent variable for denoising, but this may blur smaller texts. To preserve structure, we boost attention scores Hertz et al. (2022) for texts and preceding quotation marks, capping enhancement at 2$\times$ to prevent overamplification. To ensure control, we employ the $\tanh$ function. The $k$ token sequence is denoted as $F=\{f_{1},\dots,f_{k}\}$. The enhancement ratio is:
$\text{ratio}=1+\tanh(0.5\cdot k).$ (4)
We use a smaller ratio for single-word cases. Unlike UNet’s cross-attention, MM-DiT employs a full attention matrix with four regions: image-to-image, prompt-to-prompt, prompt-to-image, and image-to-prompt Cai et al. (2024) . We re-weight the image-to-text matrix, which controls image modality and enhances focus on visual text. At step $t$, the operation for each element in $M_{t}$ is expressed as:
$\text{Focus}(M_{t})_{i,j}=\begin{cases}\text{ratio}\cdot(M_{t})_{i,j}&\text{if}j\in F\\ (M_{t})_{i,j}&\text{otherwise.}\end{cases}$ (5)
The overall algorithmic process is illustrated in Algorithm 1.
Step 1 Step 2 Step 3 Word Accuracy↑ NED↑
$\checkmark$ 0.4422 0.6495
$\checkmark$ 0.6116 0.7750
$\checkmark$ 0.6351 0.8055
$\checkmark$ $\checkmark$ 0.6121 0.7762
$\checkmark$ $\checkmark$ 0.7328 0.8651
$\checkmark$ $\checkmark$ $\checkmark$ 0.7370 0.8679
Table 3: Ablation on CVTG-2K of Instance Fusion  (Step 1), Region Insulation  (Step 2), and Text Focus  (Step 3).
Figure 5: Qualitative comparison of TextCrafter with other baselines on CVTG-2K. TextCrafter excels in delivering harmonious and aesthetically pleasing images. It also accurately renders multiple visual texts while maintaining stability in complex scenarios.
4 Experiments
4.1 Benchmark
Currently, there is a lack of publicly available benchmark datasets specifically designed for complex visual text generation tasks. Therefore, we propose CVTG-2K, a challenging public benchmark dataset for complex visual text. All prompts are generated through the OpenAI’s O1-mini API OpenAI (2024). These prompts encompass a variety of scenes containing complex visual texts, including but not limited to street views, book covers, advertisements, posters, notes, memes, logos, and movie frames. Unlike previous standard datasets synthesized using fixed rules, CVTG-2K ensures the diversity and rationality of the data distribution. We first prompt the O1 model to conceive a scene and then imagine the spossible visual texts that may appear in that scene, utilizing Chain-of-Thought (CoT) Wei et al. (2022) techniques to ensure the quality of the generated prompts. For further details on the dialogue template with O1, refer to the supplementary material. Through rigorous filtering rules and meticulous post-processing, we have generated 2,000 prompts containing complex visual texts. On average, the visual texts in CVTG-2K contain 8.10 words and 39.47 characters, surpassing all previously published visual text benchmark datasets in terms of visual text length. Table 1 presents a comparison of the data in CVTG-2K with currently available benchmark datasets. Furthermore, CVTG-2K is the first benchmark dataset to include the number of multiple visual text regions, with the number of regions ranging from 2 to 5. We have assigned different proportions to these region numbers, approximately 20%, 30%, 30%, and 20%. With its diverse region numbers and visual text lengths, CVTG-2K provides a comprehensive evaluation of model performance in complex visual text generation tasks. Through careful human review and filtering, CVTG-2K ensures that it does not contain any discriminatory or inflammatory content.
To further enhance the challenge of CVTG-2K, we randomly added attributes to each visual text in half of the data. The attributes include size, color, and font. The sizes are categorized as large, medium, and small, and the fonts include regular, bold, italic, and cursive. All attributes are labeled using natural language to ensure that the text encoder can process them without the need for special designs. Each visual text will randomly have one or more of these attributes. This design will promote further research into the stylization and customization of visual texts in future studies. Additionally, we provide more fine-grained information. We decouple the prompts using O1-mini, separating descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$ containing multiple visual texts and expressing the relationship between each visual text and its position through the most critical words, which is the carrier of the visual text. Through meticulous manual review, we ensured the accuracy of this fine-grained information. CVTG-2K will be publicly released alongside our code, and the dataset will be available for any users adhering to the usage guidelines. For more examples of CVTG-2K, please refer to the supplementary material.
4.2 Evaluation Metrics and Baselines
The quality of CVTG are evaluated from three metrics:
  1. 1.
    Word Accuracy: The generated image is input into an OCR tool for prediction. Here, we use the SOTA open-source OCR model PPOCR-v4 PaddlePaddle (2023).
  2. 2.
    Normalized Edit Distance (NED): NED Tuo et al. (2024) is a more lenient metric than word accuracy, used to measure character-level similarity.
  3. 3.
    CLIPScore: We introduce CLIPScore, which is used to evaluate the model’s ability to follow the text prompt during the generation process Hessel et al. (2021).
We compare our method on CVTG-2K with several state-of-the-art models for visual text generation, including Stable Diffusion 3.5 Large Esser et al. (2024), FLUX.1 dev, AnyText Tuo et al. (2024) and TextDiffuser2 Chen et al. (2025). To highlight the difference between CVTG and multi-instance generation tasks, we also compare with methods like RAG-Diffusion Chen et al. (2024b) and 3DIS Zhou et al. (2025). All experiments use official or author-provided codes and configurations for optimal baseline performance. The parameter $\lambda$ described in Section 3.2 uses 0.4, and $r=5$ from Section 3.3. Notably, TextCrafter is built on FLUX.1 dev and requires no training or fine-tuning, enabling all experiments to run on a single A6000 GPU.
Figure 6: Qualitative results of ablation studies on each step individually. The incorrectly rendered visual text is highlighted with a red box, while the corresponding correctly rendered visual text is highlighted with a blue box.
4.3 Quantitative Results
As shown in Table 2, quantitative experiments on the CVTG-2K demonstrate that TextCrafter outperforms competing methods in both OCR accuracy (Word Accuracy and NED) and prompt-following ability (CLIPScore). Specifically, TextCrafter improves OCR accuracy by over 45% compared to FLUX. While general image generation models like Stable Diffusion 3.5 and FLUX perform well in simple visual text scenarios, their performance significantly degrades as the number of visual text regions increases. Methods like AnyText and TextDiffuser-2, trained on rule-based data and focusing on single-region generation, fail to generalize well to complex visual text tasks. AnyText performs better when the mask covers a larger portion of the image, leading to a slight improvement in accuracy as the number of regions increases. However, its overall performance remains 55% behind that of TextCrafter. Multi-instance control methods such as RAG-Diffusion and 3DIS, although capable of rendering complex scenes, struggle with the specific requirements of visual text generation, resulting in poor performance as complexity rises. Detailed quantitative results are provided in the supplementary material. Overall, the experiments confirm that TextCrafter excels in robustness, maintaining high accuracy and prompt-following performance even in complex visual text scenarios.
4.4 Qualitative Results
Figure 5 presents visual results comparing TextCrafter with several state-of-the-art text-to-image generation models, including SD3.5, FLUX, AnyText, TextDiffuser-2, RAG-Diffusion and 3DIS, on the CVTG-2K. While SD3.5 and FLUX produce high-quality images and can generate some visual texts, they struggle with omissions and confusion as the number of regions increases. AnyText performs suboptimally when the text consists of two or more words. TextDiffuser sacrifices background information, leading to a lower CLIPScore. RAG-Diffusion suffers from merging issues during the refinement phase. 3DIS weakens visual text information due to its layout-to-depth conversion. In contrast, TextCrafter maintains high image harmony and accurately renders multiple visual texts, adhering to prompt specifications while avoiding attribute confusion. More qualitative results are available in the supplementary material.
4.5 Ablation Study
Table 3 presents quantitative ablation results of each step individually or in combinations.
Instance Fusion.
Disabling instance fusion has a small impact on overall metrics, as it primarily establishes the correlation between visual text and its location, which doesn’t directly affect accuracy. However, without instance fusion, visual text often appears in incorrect locations. For instance, in the first column of the Figure 6, disabling instance fusion causes the text “Hot Dogs” to appear in the wrong area. Enabling instance fusion also suppresses hallucinated text, ensuring correct placement and preventing irrelevant text from appearing, as shown in the second column of the Figure 6.
Region Insulation.
Enabling region insulation improves performance across all metrics by decoupling complex visual text, reducing CVTG complexity and maximizing the pre-trained model’s ability. Region insulation achieves over 60% word accuracy on its own and reduces interference between texts. In Figure 6, region insulation prevents interference, improving visual text clarity and accuracy, as seen in the corrected rendering of “Arrival” in the fourth column.
Text Focus.
Text focus has the most significant impact on improving metrics, especially for small-size visual text. It enhances both clarity and accuracy, as seen in the clear generation of “Next Stop” on the train in the fifth column in Figure 6, and “VIP Access” in the last column. When enabled alone, text focus achieves 63.51% word accuracy but is about 10% lower than the optimal performance, indicating that it requires support from instance fusion and region insulation for stability, particularly in complex visual text scenarios.
5 Conclusion
This paper delves into the complex visual text generation (CVTG) task, which involves generating textual content across multiple regions of an image. We propose TextCrafter, a novel training-free framework based on DiT framework, employing step-by-step refinement from prompts to clauses, and finally to specific visual text for precise generation. Key innovations in TextCrafter include instance fusion to link text with its spatial carrier, the region insulation strategy to reduce complexity and prevent text loss, and a text focus approach for accurate small-text rendering. We also construct a new CVTG-2K benchmark dataset for the CVTG task. Extensive experiments demonstrate TextCrafter’s clear advantages against the existing competitors.
Ethical Statement
Any users of our model are prohibited from using it to create false or malicious information.