### v3

# TextCrafter: Accurately Rendering Multiple Texts in Complex Visual Scenes

Abstract
This paper explores the task of Complex Visual Text Generation (CVTG), which centers on generating intricate textual content distributed across diverse regions within visual images. In CVTG, image generation models often rendering distorted and blurred visual text or missing some visual text. To tackle these challenges, we propose TextCrafter, a novel multi-visual text rendering method. TextCrafter employs a progressive strategy to decompose complex visual text into distinct components while ensuring robust alignment between textual content and its visual carrier. Additionally, it incorporates a token focus enhancement mechanism to amplify the prominence of visual text during the generation process. TextCrafter effectively addresses key challenges in CVTG tasks, such as text confusion, omissions, and blurriness. Moreover, we present a new benchmark dataset, CVTG-2K, tailored to rigorously evaluate the performance of generative models on CVTG tasks. Extensive experiments demonstrate that our method surpasses state-of-the-art approaches.
Figure 1: TextCrafter enables precise multi-region visual text rendering, addressing the challenges of long, small-size, various numbers, symbols and styles in visual text generation. We illustrate the comparisons among TextCrafter with three state-of-the-art models, i.e., FLUX BlackForest [1], TextDiffuser-2 Chen et al. [5] and 3DIS Zhou et al. [45].
1 Introduction
In recent years, diffusion models  [13, 27, 30, 28, 29, 40, 8, 6, 17] have emerged as the forefront technology in visual image generation. Recent studies, represented by general-purpose text-to-image models such as Flux [1] and SD 3.5 [9], have acquired the ability to render simple visual text through large-scale pretraining. However, when confronted with complex real-world visual text scenarios, these models often struggle with critical challenges such as text distortion, omission, and blurriness. These shortcomings significantly hinder their practical applicability in real-world use cases. Achieving precise, legible, and contextually accurate visual text rendering in complex scenes remains essential yet highly challenging.
Recent studies in multi-instance control [10, 43, 7, 45, 18] primarily focus on generating multiple instances within a single image simultaneously, striving to achieve precise regulation over factors such as quantity, position and attributes. However, a crucial yet often overlooked category in the visual domain is visual text. Text is an indispensable component of the real world, distinguished by its unparalleled complexity and delicacy compared to conventional objects, flora, or fauna. Unlike these entities, text possesses an exceptionally intricate internal structure, where even minor alterations can cause significant changes in visual appearance, potentially leading to misrecognition or complete illegibility. Existing multi-instance approaches, such as MIGC [43] and 3DIS [45], primarily focus on sentence-level information and interaction, offering a limited understanding of text at a finer granularity. Moreover, RPG [35] and RAG-Diffusion [7] adjust instance structures during attention fusion but fall short of the precision and stability needed for visual text rendering.
Recent studies on visual text rendering primarily focus on accurately generating a single text region. For instance, methods like AnyText [32], Glyph-byT5 [23], SceneTextGen [39], Diff-Text [38], and TextDiffuser [4] often employ specialized modules to encode visual text representations and condition the diffusion model accordingly. However, these methods face two major limitations: 1) They rely on predefined rules to synthesize visual text training datasets. Due to the need for highly accurate text annotations, manual verification is often required, making the creation of complex visual text datasets both labor-intensive and costly. 2) Most approaches depend on fine-tuned text encoders (e.g., Glyph-by-T5) or trained conditional control encoders (e.g., AnyText, Diff-Text, TextDiffuser, SceneTextGen) to facilitate text generation. In complex multi-text generation, however, these approaches often cause interference among the control features of different targets, making it difficult to balance fine-grained local control with global consistency across multiple texts.
This paper explores the task of Complex Visual Text Generation (CVTG), which aims to generate intricate text content distributed across different regions. To tackle the challenges of CVTG, we propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes. TextCrafter enhances multi-text control, ensuring accurate text presentation and logical layout. Our framework comprises three key stages: 1) Instance Fusion: This stage strengthens the connection between visual text and its corresponding carrier, ensuring strong consistency between the text content and its surrounding environment. It effectively prevents text from appearing in incorrect positions. 2) Region Insulation: Leveraging the positional priors of the pre-trained DiT model [26] (e.g., FLUX), this step initializes the layout information for each text instance while separating and denoising text prompts across different regions. This process prevents early interference between text areas, reducing confusion and the risk of content omission in multi-text scenarios. 3) Text Focus: An attention control mechanism is introduced to enhance the attention maps of visual text, refining the fidelity of text rendering. This stage ensures textual accuracy and addresses issues of blurriness, particularly in smaller text. We present some examples in Figure 1. Our contributions are summarized as follows:
  • We propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes, consisting of three key stages: instance fusion, region insulation, and token focus.
  • We construct a new CVTG-2K dataset, which encompasses a diverse range of visual text prompts varying in position, quantity, length, and attributes, serving as a robust benchmark for evaluating the performance of generation models in CVTG tasks.
  • We conduct extensive quantitative and qualitative experiments on the CVTG-2K benchmark. The results demonstrate that TextCrafter delivers exceptional performance, proving its superior effectiveness and robustness in tackling the CVTG task.
Figure 2: Overview of our TextCrafter. TextCrafter consists of three steps. (a) Instance Fusion: strengthen the connection between visual text and its corresponding carrier. (b) Region Insulation: leverage the positional priors of the pre-trained DiT model to initialize the layout information for each text instance while separating and denoising text prompts across different regions. (c) Text Focus: enhance the attention maps of visual text, refining the fidelity of text rendering.
2 Related Works
Multi-instance Compositional Generation.
Attend-and-Excite [3] and BOX-Diffusion [34] guide diffusion models to align with prompts. GLIGEN [21], MIGC [43], and CreativeLayout [37] incorporate bounding boxes into trainable layers. RPG [35], RAG-Diffusion [7] and DreamRenderer [44] decompose generation into regional tasks. However, these methods overlook visual text, a critical element for image realism.
Visual Text Generation.
DiffSTE [19], Glyph-ByT5 [23], and UDiffText [41] use character-level encoders for word appearance. TextDiffuser [4] and GlyphControl [36] generate text layout masks. Diff-Text [38] restricts attention maps. GlyphDraw [24] utilizes dual models for positioning and rendering. TextDiffuser-2 [5] leverages LLMs for layout planning. These approaches optimize for single region or individual object, limiting generalization in complex scenes.
Benchmark for Visual Text.
GlyphControl introduced SimpleBench and CreativeBench [36], but these contain only single-word scenes. MARIO-Eval [4] has low-quality prompts. DrawTextExt Creative [24] lacks sufficient coverage. AnyText-benchmark [32] treats words separately, limiting distribution. Existing benchmarks focus on a single object or location, failing to capture real-world complexity. Thus, we propose CVTG-2K to evaluate models with diverse positions, quantities, lengths, and attributes in complex visual text scenarios.
Figure 3: Illustration of tokenizing the prompt “A sidewalk poster with ‘Register Now for IJCAI 2025’.” along with the attention map corresponding to each token. The use of preceding quotation marks can reinforce the relationship between text tokens and carrier tokens.
Figure 4: Motivation for using a layout optimizer. LLM/MLLM-based layout designers frequently generate overlapping bounding boxes, which do not align with the layout preferences of Diffusion models. Our proposed layout optimizer addresses this limitation without requiring additional fine-tuning.
3 Methodology
3.1 Overview
Problem Definition.
In CVTG, users provide a global prompt $P$ containing multiple visual text descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$, where each description includes the text’s main body and descriptors (position, attributes). The main text bodies are defined as $VT=\{vt_{1},vt_{2},\dots,vt_{n}\}$, with each $vt_{i}$ corresponding to descriptors $d_{i}$. The model must generate an image from prompt $P$ where each $vt_{i}$ appears with its corresponding description $d_{i}$.
Challenges in CVTG.
Complex prompts cause interference among visual texts, leading to three degradation issues: $1$) Text confusion: Visual texts $vt_{i}$ and $vt_{j}$ intertwine, generating erroneous characters/words. $2$) Text omission: Only text from description $d_{i}$ appears, neglecting $vt_{j}$ from $d_{j}$. $3$) Text blurriness: Small-scale text in $d_{i}$ receives insufficient attention, appearing blurred in the generated image.
Motivation.
Current text-to-image models excel with individual visual texts. By decomposing CVTG into simpler sub-tasks, overall complexity decreases. Solving these sub-tasks separately and combining their solutions accomplishes the CVTG task. This approach shifts focus progressively from the entire task to sub-tasks then to specific visual texts, enabling precise rendering. Based on this coarse-to-fine strategy, we present TextCrafter (Figure 2), detailed in subsequent sections.
3.2 Instance Fusion
It is crucial to prevent visual texts from floating in the image and ensure proper positioning. We propose an instance fusion strategy that integrates the visual text’s content with its spatial carrier. Rather than using the text’s own embedding, we leverage embedding of the preceding quotation mark to establish spatial relationships. As illustrated in Figure 3, the preceding quotation mark in the attention map corresponds to the visual text it governs, indicating that it encapsulates complete information about the text. According to token additivity [14], we incorporate the quotation mark’s embedding into the carrier’s embedding via weighted fusion to ensure precise alignment, where $\lambda$ controls the proportion of the addition.
Figure 5: For a pre-trained DiT model, only a few denoising steps are required to approximate the image layout and subject positions. After 8 steps, the layout closely resembles that of a full 50-step process, with subsequent iterations primarily refining details.
Table 1: Comparison of CVTG-2K with existing public visual text Benchmarks in terms of Number of Samples (Number), Average Word Count (Word), Average Character Count (Char), Attribute, and number of Regions (Region).
Benchmark Number Word Char Attribute Region
CreativeBench 400 1.00 7.29 $\times$ $\times$
MARIOEval 5414 2.92 15.47 $\times$ $\times$
DrawTextExt 220 3.75 17.01 $\times$ $\times$
AnyText-benchmark 1000 4.18 21.84 $\times$ $\times$
CVTG-2K 2000 8.10 39.47
$\checkmark$
size/color/font
$\checkmark$
2/3/4/5
Algorithm 1 TextCrafter Image Generation
1:  Input: The encoded embedding of a global prompt $P$ is denoted as $c$. The encoded embedding of multiple descriptions $D$ is denoted as $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. 2:  Parameter: the proportion of the addition $\lambda$, the $r$ steps of Region Insulation, the pre-generation steps $\tau$. 3:  Output: A generated latent $x_{0}$ 4:  {Step 1: Instance Fusion} 5:  $c=\text{InstanceFusion}(c,\lambda)$ 6:  {Step 2: Region Insulation} 7:  $z_{t}\sim\mathcal{N}(0,1)$ 8:  for $t=\tau,\tau-1,\dots,1$ do 9:     $z_{t-1},A_{t}\leftarrow\text{DiT}(z_{t},c,t)$ 10:  end for 11:  for $i=1,2,\dots,n$ do 12:     $p_{i}\leftarrow(0,0)$ 13:     $p_{\text{max},i}\leftarrow\arg\max_{p_{i}}A_{t}(p_{i},i)$ 14:  end for 15:  $bbx_{1},\dots,bbx_{n}\leftarrow\text{MILP}(p_{\text{max},1},\dots,p_{\text{max},n})$ 16:  $x_{T}\sim\mathcal{N}(0,1)$ 17:  for $i=1,2,\dots,n$ do 18:     $\hat{x}_{t}^{i}\leftarrow\text{Initialize}(\mathcal{N}(0,1),bbx_{i})$ 19:  end for 20:  for $t=T,T-1,\dots,T-r+1$ do 21:     $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$ 22:     for $i=1,2,\dots,n$ do 23:        $\hat{x}_{t-1}^{i}\leftarrow\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$ 24:        $x_{t-1}\leftarrow\text{Reinsert}(x_{t-1},\hat{x}_{t-1}^{i},bbx_{i})$ 25:     end for 26:  end for 27:  {Step 3: Text Focus} 28:  for $t=T-r,T-r-1,\dots,1$ do 29:     for $\text{AttentionLayer}_{i}\in\epsilon_{\theta}$ do 30:        $M_{t}=Q_{t}^{i}\times K_{t}^{i}$ 31:        $M_{t}\leftarrow\text{Focus}(M_{t})$ 32:        $\text{Attention}_{t}^{i}\leftarrow\text{softmax}(M_{t}/\sqrt{d})\times V_{i}$ 33:     end for 34:     $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$ 35:  end for 36:  return  $x_{0}$
Table 2: Quantitative results on the proposed CVTG-2K. The results highlight TextCrafter’s superiority across five metrics.
Model Word Accuracy↑ NED↑ CLIPScore↑ VQAScore↑ Aesthetics↑
FLUX.1 dev (2024) 0.4965 0.6879 0.7401 0.8886 5.91
SD3.5 Large (ICML 2024) 0.6548 0.8470 0.7797 0.9297 5.56
AnyText (ICLR 2024) 0.1804 0.4675 0.7432 0.6935 4.53
TextDiffuser-2 (ECCV 2024) 0.2326 0.4353 0.6765 0.5627 4.51
3DIS (ICLR 2025) 0.3813 0.6505 0.7767 0.8684 4.86
TextCrafter (FLUX) 0.7370 0.8679 0.7868 0.9140 5.70
TextCrafter (SD3.5) 0.7600 0.9038 0.8023 0.9299 5.45
3.3 Region Insulation
Pre-generation.
In CVTG, multiple texts often influence each other’s glyphs and layout [42]. To simplify, we decouple them into independent instances. Specifically, we assign a rectangular bounding box to each visual text to represent its layout, initialized based on the positional preferences of the pre-trained DiT model. This allows us to effectively isolate the texts, preventing interference during generation. Instead of manually specifying layouts or relying on LLMs, we leverage the diffusion model’s learned positional knowledge [31, 16]. Figure 4 demonstrates a notable limitation of LLMs/MLLMs: despite dedicated fine-tuning, the layouts they provide often exhibit overlapping bounding boxes and do not conform to the layout preferences inherent in Diffusion models. As shown in Figure 5, during the early denoising steps, we capture attention maps and identify the points of maximum attention for each visual text. This pre-generation process guides layout initialization. Assume that the attention map obtained during this process is $A_{t}$, where $t$ represents the current step, and $p$ stands for pixel:
$p_{\text{max}}=\arg\max_{p}A_{t}(p).$ (1)
Layout Optimizer.
Next, the rectangular bounding boxes for each visual text are optimized using Mixed-Integer Linear Programming (MILP) to minimize the Manhattan distance between the bounding box center and the attention point. Let $p_{\text{max},i}$ be the maximum attention point of $vt_{i}$ and $c_{i}$ the rectangular bounding box center. The optimization objective is:
$\min\sum_{i}D_{\text{Manhattan}}(p_{\text{max},i},c_{i}).$ (2)
Simultaneously incorporate a set of constraints to ensure that the layout of each visual text is reasonable and non-overlapping. The solution to this optimization problem can be represented as assigning a bounding box $bbx_{i}$ $(m_{\text{offset}},n_{\text{offset}},m_{\text{scale}},n_{\text{scale}})$ to each description. Then, we encode the prompt $P$ and the corresponding descriptions of visual text $D=\{d_{1},d_{2},\dots,d_{n}\}$, obtaining values for $c$ and $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. $c$ needs to be processed through instance fusion introduced in Section 3.2. The latent $\hat{x}_{i}$ is conditioned on $\hat{c}_{i}$, and the initial latent $x$ is conditioned on the complete $c$. When $t>T-r$ (i.e., during the first $r$ steps of the denoising process), denoising follows the formula:
$x_{t-1}=x_{t}-\epsilon_{\theta}(x_{t},c),\quad\hat{x}_{t-1}^{i}=\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$ (3)
where $\epsilon_{\theta}$ represents the noise predictor. After each denoising step, $\hat{x}_{t}^{i}$ is reinserted into the corresponding region in $x_{t}$.
3.4 Text Focus
Region insulation creates distinct visual boundaries between visual texts. To enhance harmony, regions are merged into a single latent variable for denoising, but this may blur smaller texts. To preserve structure, we boost attention scores [11, 15] for texts and preceding quotation marks, capping enhancement at 2$\times$ to prevent overamplification. To ensure control, we employ the $\tanh$ function. The $k$ token sequence is denoted as $F=\{f_{1},\dots,f_{k}\}$. The enhancement ratio is:
$\text{ratio}=1+\tanh(0.5\cdot k).$ (4)
We use a smaller ratio for single-word cases. Unlike UNet’s cross-attention, MM-DiT employs a full attention matrix with four regions: image-to-image, prompt-to-prompt, prompt-to-image, and image-to-prompt [2] . We re-weight the image-to-text matrix, which controls image modality and enhances focus on visual text. At step $t$, the operation for each element in $M_{t}$ is expressed as:
$\text{Focus}(M_{t})_{i,j}=\begin{cases}\text{ratio}\cdot(M_{t})_{i,j}&\text{if}j\in F\\ (M_{t})_{i,j}&\text{otherwise.}\end{cases}$ (5)
The overall algorithmic process is illustrated in Algorithm 1.
Table 3: Ablation studies on CVTG-2K dataset, including Instance Fusion (Step 1), Region Insulation (Step 2), Text Focus (Step 3), and Layout methods.
Steps Layout Method
Metrics S1 S2 S3 S1+S2 S2+S3 All Random Fixed Optimizer
Word Accuracy↑ 0.4422 0.6116 0.6351 0.6121 0.7328 0.7370 0.6900 0.6994 0.7370
NED↑ 0.6495 0.7750 0.8055 0.7762 0.8651 0.8679 0.8374 0.8475 0.8679
Figure 6: Qualitative comparison of TextCrafter with other baselines on CVTG-2K. TextCrafter excels in delivering harmonious and aesthetically pleasing images. It also accurately renders multiple visual texts while maintaining stability in complex scenarios.
4 Experiments
4.1 Benchmark
Currently, there is a lack of publicly available benchmark datasets specifically designed for complex visual text generation tasks. Therefore, we propose CVTG-2K, a challenging public benchmark dataset for complex visual text. All prompts are generated through the OpenAI’s O1-mini API [25]. These prompts encompass a variety of scenes containing complex visual texts, including but not limited to street views, book covers, advertisements, posters, notes, memes, logos, and movie frames. Unlike previous standard datasets synthesized using fixed rules, CVTG-2K ensures the diversity and rationality of the data distribution. We first prompt the O1 model to conceive a scene and then imagine the spossible visual texts that may appear in that scene, utilizing Chain-of-Thought (CoT) [33] techniques to ensure the quality of the generated prompts. For further details on the dialogue template with O1, refer to the supplementary material. Through rigorous filtering rules and meticulous post-processing, we have generated 2,000 prompts containing complex visual texts. On average, the visual texts in CVTG-2K contain 8.10 words and 39.47 characters, surpassing all previously published visual text benchmark datasets in terms of visual text length. Table 1 presents a comparison of the data in CVTG-2K with currently available benchmark datasets. Furthermore, CVTG-2K is the first benchmark dataset to include the number of multiple visual text regions, with the number of regions ranging from 2 to 5. We have assigned different proportions to these region numbers, approximately 20%, 30%, 30%, and 20%. With its diverse region numbers and visual text lengths, CVTG-2K provides a comprehensive evaluation of model performance in complex visual text generation tasks. Through careful human review and filtering, CVTG-2K ensures that it does not contain any discriminatory or inflammatory content.
To further enhance the challenge of CVTG-2K, we randomly added attributes to each visual text in half of the data. The attributes include size, color, and font. The sizes are categorized as large, medium, and small, and the fonts include regular, bold, italic, and cursive. All attributes are labeled using natural language to ensure that the text encoder can process them without the need for special designs. Each visual text will randomly have one or more of these attributes. This design will promote further research into the stylization and customization of visual texts in future studies. Additionally, we provide more fine-grained information. We decouple the prompts using O1-mini, separating descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$ containing multiple visual texts and expressing the relationship between each visual text and its position through the most critical words, which is the carrier of the visual text. Through meticulous manual review, we ensured the accuracy of this fine-grained information. CVTG-2K has been publicly released alongside our code, and the dataset is available for any users adhering to the usage guidelines. For more examples of CVTG-2K, please refer to the supplementary material.
4.2 Evaluation Metrics and Baselines
The quality of CVTG are evaluated from five metrics: Word Accuracy and Normalized Edit Distance (NED) [32] are employed to evaluate the accuracy of text rendering. CLIPScore [12], VQAScore [22], and Aesthetics [20] are utilized to assess the quality of the generated images. For detailed configurations of the metrics calculation, please refer to the supplementary materials.
We compare our method on CVTG-2K with several state-of-the-art Text-to-Image models including Stable Diffusion 3.5 Large [9], FLUX.1 dev. To highlight the difference between CVTG and multi-instance generation tasks, we also compare with methods like 3DIS [45]. To highlight the distinction between CVTG and multi-line text generation tasks, we conducted comparative analyses of methods capable of generating multi-line text, such as AnyText [32] and TextDiffuser2 [5]. The parameter $\lambda$ described in Section 3.2 uses 0.4, and $r=5$ from Section 3.3. Notably, TextCrafter requires no training or fine-tuning and is compatible with different DiT backbone networks. We conducted experiments using both FLUX.1 dev and Stable Diffusion 3.5 Large as backbone networks. All experiments can be completed on a single A6000 GPU. For specific data on memory utilization and computational runtime, please refer to the supplementary materials.
Figure 7: Qualitative results of ablation studies on each step individually. The incorrectly rendered visual text is highlighted with a red box, while the corresponding correctly rendered visual text is highlighted with a blue box.
4.3 Quantitative Results
As demonstrated in Table 2, TextCrafter surpasses competing methods in both text accuracy and image quality on CVTG-2K, with TextCrafter(FLUX) improving word accuracy by over 45% relative to FLUX. While Stable Diffusion 3.5 and FLUX perform adequately in simple scenarios, their efficacy diminishes significantly with increased textual complexity. AnyText and TextDiffuser-2, despite being trained on rule-based data, fail to generalize to multi-region tasks, with AnyText performing 55% below TextCrafter even with improved mask coverage. Similarly, 3DIS struggles with visual text generation despite its multi-instance capabilities. As a training-free approach, TextCrafter maintains superior aesthetic scores compared to specialized alternatives while balancing textual accuracy with visual coherence, demonstrating exceptional robustness across complexity levels. Comprehensive results are available in the supplementary material.
4.4 Qualitative Results
Figure 6 illustrates comparative visual results between TextCrafter and state-of-the-art models (SD3.5, FLUX, AnyText, TextDiffuser-2, and 3DIS) on CVTG-2K. While SD3.5 and FLUX generate visually appealing images, they exhibit deficiencies in text rendering as regional complexity increases. AnyText demonstrates suboptimal performance with multi-word text instances, TextDiffuser-2 compromises background fidelity (resulting in diminished CLIPScore), and 3DIS deteriorates textual information during layout-to-depth conversion. Conversely, TextCrafter achieves superior visual harmony while accurately rendering multiple text regions without attribute confusion or prompt deviation. Additional qualitative evaluations are presented in the supplementary material.
4.5 Ablation Study
Table 3 presents quantitative ablation results for individual and combined components.
Instance Fusion.
Disabling instance fusion minimally impacts overall metrics but is vital for correct visual text localization; its absence often leads to misplaced text (e.g., "Hot Dogs," Figure 7, first column). Enabling it also suppresses hallucinated text, ensuring accurate placement (Figure 7, second column).
Region Insulation.
Enabling region insulation improves all metrics by decoupling complex visual text, thereby reducing CVTG complexity and maximizing pre-trained model utility. It achieves over 60% word accuracy independently and reduces inter-text interference, enhancing clarity (e.g., "Arrival," Figure 7, fourth column).
Text Focus.
Text focus offers the most significant metric improvement, especially for small visual text, by enhancing clarity and accuracy (e.g., "Next Stop" and "VIP Access," Figure 7, fifth and last columns). Alone, it yields 63.51% word accuracy, roughly 10% below optimal, indicating a need for instance fusion and region insulation for stability in complex scenarios.
Layout Optimizer Effectiveness.
The proposed Manhattan distance-based layout optimization significantly influences final generation by creating non-overlapping bounding boxes aligned with pretrained models’ spatial preferences. Table 3 quantitatively demonstrates its superiority over alternatives like Random or Fixed Layouts.
5 Conclusion
This paper addresses complex visual text generation (CVTG) by proposing TextCrafter, a novel training-free framework built on the DiT framework. It employs step-by-step refinement from prompts to clauses and finally to specific visual text. Key innovations include Instance Fusion for text-spatial carrier linkage, Region Insulation to reduce complexity and text loss, and a Text Focus approach for accurate small-text rendering. A new CVTG-2K benchmark dataset is also introduced. Extensive experiments demonstrate TextCrafter’s significant advantages over existing methods.
Limitations
TextCrafter has two main limitations: (1) degraded performance with conflicting requirements (e.g., large text in small areas), where balancing spatial constraints and clarity is challenging; and (2) as a training-free method, its text enhancement can introduce artifacts that affect overall image aesthetics, especially in visually dense scenes. Future work will aim to develop mechanisms for high-fidelity text rendering while maintaining the aesthetic integrity of the image.